OpenStudy (anonymous):

show the work of the function f(x)=(x-3^1/2(2x-5) that has a derivative of f'(x)=(6x-17/2(x-3)^1/2)

OpenStudy (anonymous):

\[f(x)=\frac{x-3}{2(2x-5)}\] \[f'(x)=\frac{4x-10-4x+12}{(4x-10)^2}\] \[=\frac{2}{(4x-10)^2}\]

OpenStudy (anonymous):

2x^2-5x-6x^(3/2)-15x^1/2 =4x-3/2x^1/2-5